# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 2 Group splittings and their halfspaces
The combination of (3), (4) and (6) tells us that $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry, as required. Thus the halfspaces from the second definition will be quasi-isometric to the halfspaces from the third definition. We can also deduce from this that the halfspaces from both the second and third definitions are coarsely well-defined – for instance, if $S^{\prime}$ is another finite generating set for $G$, and $\mathfrak{h}^{\prime}_{1}\subseteq\operatorname{Cay}(G,S^{\prime})$ is the halfspace with vertex set $V\mathfrak{h}^{\prime}_{1}=f^{-1}(\mathfrak{h})$, then the above argument tells us that the halfspaces $\mathfrak{h}_{1}$ and $\mathfrak{h}^{\prime}_{1}$ are both quasi-isometric to $\mathfrak{h}_{2}$, hence they are quasi-isometric to each other. ∎
@@
As in Definition 2, let $p:X\to T$ be a tree of spaces for $G\curvearrowright T$. We must show that $p^{-1}(\mathfrak{h})$ is one-ended, so suppose for contradiction that $K\subseteq p^{-1}(\mathfrak{h})$ is a finite subgraph such that $p^{-1}(\mathfrak{h})$ has multiple unbounded components. We are assuming that $G$ and $G_{v}$ are one-ended, hence $X$ and $p^{-1}(v)$ are one-ended as well. In particular $p^{-1}(v)-K$ only has one unbounded component, call it $C_{1}$. Now consider the finite subgraph
@@
Let $C_{2}$ be an unbounded component of $p^{-1}(\mathfrak{h})-K$ disjoint from $C_{1}$. Observe that $p^{-1}(\hat{e})$ is contained in a bounded neighborhood of $p^{-1}(v)$ with respect to the induced metric of $p^{-1}(\mathfrak{h})$, hence also in a bounded neighborhood of $C_{1}$. It follows that $C_{2}\cap p^{-1}(\hat{e})$ is bounded, so only finitely many vertices in $C_{2}$ lie in $p^{-1}(\hat{e}\cup v)$. Any vertex in $C_{2}-p^{-1}(\hat{e}\cup v)$ is separated from $C_{1}$ in $X$ by $K^{\prime}$, so this contradicts one-endedness of $X$. ∎
@@ 3 Pocsets
Elements $A,B\in\mathcal{P}$ are nested if $A\leq B$ or $B\leq A$. And $A,B\in\mathcal{P}$ are transverse if neither $A,B$ nor $A,B^{*}$ are nested. The width of $\mathcal{P}$ is the maximum number of pairwise transverse elements, if such a maximum exists, otherwise we say the width is $\infty$.
@@
An ultrafilter $\omega$ is DCC (descending chain condition) if it contains no strictly descending infinite chain $A_{1}>A_{2}>A_{3}>\cdots$.
@@
Let $\mathcal{P}$ be a pocset of finite width that admits at least one DCC ultrafilter. Then there is a CAT(0) cube complex $C=C(\mathcal{P})$, called the cubing of $\mathcal{P}$, such that:
@@
In this paper we will only deal with pocsets whose cubings are trees, or equivalently, pocsets that contain no pairs of transverse elements. One source of pocsets will come from the following definition.
@@
The action $G\curvearrowright T$ is minimal, so $p(\mathfrak{h})$ is not contained in a bounded neighborhood of $p(\hat{\mathfrak{h}})$ or $p(\mathfrak{h}^{*})$. Since $p$ is distance non-increasing, we have that $\mathfrak{h}$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}$ or $\mathfrak{h}^{*}$. Secondly, as $\hat{\mathfrak{h}}$ separates $\mathfrak{h}-\hat{\mathfrak{h}}$ from $\mathfrak{h}^{*}-\hat{\mathfrak{h}}$ in $X$, and since $X$ is one-ended, we must have that $\hat{\mathfrak{h}}$ is unbounded. ∎
@@ 4.1 Minimal cuts
Definition 12. Let $Y$ be a connected simplicial graph. For $C\subseteq VY$, define the edge boundary of $C$, denoted $\delta C$, to be the set of edges in $Y$ with one vertex in $C$ and one vertex in $VY-C$. A ray in $Y$ is an embedded one-way infinite path in $Y$. A cut is a subset $C\subseteq VY$ with $\delta C$ finite and such that the induced subgraphs of $C$ and $VY-C$ are both connected and contain rays. Note that, if $Y$ is locally finite, then a connected subgraph of $Y$ contains a ray if and only if it is unbounded. A cut $C$ is minimal if $|\delta C|$ is minimal.
@@
Clearly, $\mathcal{P}(Y,C)$ is a family of non-empty subsets of $VY$ that is closed under complementation. It remains to show that, for $x,y\in VY$, the set $S=\{A\in\mathcal{P}(Y,C)\mid x\in A,\,y\notin A\}$ is finite. Indeed, fix a path $\gamma$ from $x$ to $y$. Given $A\in S$, we either have $A=gC$ or $A=VY-gC$ for some $g\in\operatorname{Aut}(Y)$. Now, $\delta(gC)$ determines the pair $(gC,VY-gC)$, so it suffices to show that there are only finitely many possibilities for $\delta(gC)$. But the path $\gamma$ contains an edge in $\delta(gC)$, and $\delta(gC)$ is an $\operatorname{Aut}(Y)$-translate of the finite edge set $\delta C$, so the local finiteness of $Y$ implies that there are only finitely many possibilities for $\delta(gC)$. ∎
@@
4.2 Chopping up a halfspace $\mathfrak{h}_{0}$
We now turn to the proof of Theorem 4.1, which will be spread over Subsections 4.24.5. Let $G\curvearrowright T$ be a splitting as in the theorem, and suppose that it does not have one-ended halfspaces. Let $p:X\to T$ be a tree of spaces associated to $G\curvearrowright T$ (Definition 2).
Define a halfspace cut to be a pair $(\mathfrak{h},C)$ where $\mathfrak{h}\in\mathcal{H}(X)$ and $C$ is a cut in $\mathfrak{h}$. As in Definition 4.1, let $\delta C$ denote the set of edges in $\mathfrak{h}$ with one vertex in $C$ and one vertex in $V\mathfrak{h}-C$. As the splitting $G\curvearrowright T$ does not have one-ended halfspaces, there exists a halfspace $\mathfrak{h}\in\mathcal{H}(X)$ that is not one-ended, and there exists $C\subseteq V\mathfrak{h}$ such that $(\mathfrak{h},C)$ is a halfspace cut.

### v4

@@ 2 Group splittings and their halfspaces
The combination of (6), (7) and (9) tells us that $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry, as required. ∎
Lemma 2.6.
The second and third definitions for halfspaces given above are coarsely well-defined and coarsely equivalent. More precisely, if we fix a splitting $G\curvearrowright T$ and a halfspace $\mathfrak{h}$ of $T$, then the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ remains the same up to quasi-isometry when we
  1. (1)
    switch between the two definitions of halfspace,
  2. (2)
    change the finite generating set for $G$ or the orbit map $f$ from Definition 2, or
  3. (3)
    change the graph $X$ from Definition 2.
In particular, the notion of the halfspaces of a splitting $G\curvearrowright T$ being one-ended is well-defined.
Proof.
Let $S$ be a finite generating set for $G$ as in Definition 2 (so the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ is connected for every halfspace $\mathfrak{h}\in\mathcal{H}(T)$). Let $\mathfrak{h}\in\mathcal{H}(T)$ be a halfspaces of $T$. Let $\mathfrak{h}_{1}\subset\operatorname{Cay}(G,S)$ and $\mathfrak{h}_{2}\subset X$ be the halfspaces corresponding to $\mathfrak{h}$ from Definitions 2 and 2 respectively. Lemma 2.5 implies that $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ are quasi-isometric, which proves item (1). Moreover, we can keep using Lemma 2.5 to show that $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ remain the same up to quasi-isometry when making the changes indicated in (2) and (3). For instance, if $S^{\prime}$ is another finite generating set for $G$ as in Definition 2, and $\mathfrak{h}^{\prime}_{1}\subseteq\operatorname{Cay}(G,S^{\prime})$ is the halfspace corresponding to $\mathfrak{h}$ and $S^{\prime}$ from Definition 2, then Lemma 2.5 implies that the halfspaces $\mathfrak{h}_{1}$ and $\mathfrak{h}^{\prime}_{1}$ are both quasi-isometric to $\mathfrak{h}_{2}$, hence they are quasi-isometric to each other. Similarly, the halfspace $\mathfrak{h}_{1}$ remains the same up to quasi-isometry if we change the orbit map $f$ from Definition 2, and the halfspace $\mathfrak{h}_{2}$ remains the same up to quasi-isometry if we change the graph $X$ from Definition 2. ∎
Lemma 2.7.
The first definition of halfspace is coarsely well-defined and coarsely equivalent to the second definition in the following sense.
  1. (1)
    Let $G=A*_{C}B$ be an amalgam (resp. $G=A*_{C}$ an HNN extension) as in Definition 2. Then the halfspaces $\mathfrak{h}_{A},\mathfrak{h}_{B}$ (resp. $\mathfrak{h}^{+},\mathfrak{h}^{-}$) from Definition 2 remain the same up to quasi-isometry when we change the choice of finite generating sets $S_{A},S_{B}$ (resp. $S_{A}$). Furthermore, if $G\curvearrowright T$ is the Bass-Serre tree associated to the given amalgam or HNN decomposition of $G$, then each halfspace from Definition 2 is quasi-isometric to a halfspace of $G\curvearrowright T$ from Definition 2, and vice versa.
  2. (2)
    Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix an edge $e\in ET$ and let $\mathfrak{h}\in\mathcal{H}(T)$ be one of the halfspaces associated to $e$. Then $G$ splits as either an amalgam $G=A*_{C}B$ or an HNN extension $G=A*_{C}$, with $C=G_{e}$ in either case, and so that the halfspace from Definition 2 corresponding to $\mathfrak{h}$ is quasi-isometric to one of the halfspaces $\mathfrak{h}_{A},\mathfrak{h}_{B},\mathfrak{h}^{+},\mathfrak{h}^{-}$ from Definition 2.
Proof.
  1. (1)
    We give the proof for the case where $G=A*_{C}B$ is an amalgam (the HNN case is similar). Let $G\curvearrowright T$ be the corresponding Bass–Serre tree and let $v_{A},v_{B}\in VT$ be the vertices stabilized by $A$ and $B$ respectively. Let $e$ be the edge joining $v_{A}$ and $v_{B}$ and let $\mathfrak{h}\subseteq T$ (resp. $\mathfrak{h}^{*}$) be the halfspace bounded by $\hat{e}$ that contains $v_{A}$ (resp. $v_{B}$). Let $f:G\to VT$ be the orbit map given by $g\mapsto gv_{A}$ and let $S=S_{A}\cup S_{B}$ be the finite generating set for $G$ from Definition 2. It is not hard to show that the halfspace $\mathfrak{h}_{A}$ from Definition 2 is precisely the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$. Then, by Lemma 2.5, $\mathfrak{h}_{A}$ is quasi-isometric to the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ from Definition 2. And then, by Lemma 2.6, $\mathfrak{h}_{A}$ is also quasi-isometric to the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ from Definition 2 (with respect to any choice of orbit map and any suitable finite generating set $S^{\prime}$ for $G$). An analogous argument shows that the halfspace $\mathfrak{h}_{B}$ from Definition 2 is quasi-isometric to the halfspace from Definition 2 corresponding to $\mathfrak{h}^{*}$. Of course the $G$-translate of a given halfspace (with either definition) is isometric to the given halfspace, hence every halfspace from Definition 2 is quasi-isometric to a halfspace from Definition 2, and vice versa. Finally, the halfspaces $\mathfrak{h}_{A},\mathfrak{h}_{B}$ from Definition 2 remain the same up to quasi-isometry when we change the choice of finite generating sets $S_{A},S_{B}$ because the halfspaces from Definition 2 are coarsely well-defined (Lemma 2.6).
  2. (2)
    Let $T^{\prime}$ be obtained from $T$ by collapsing all edges which are not in the orbit of $e$. Then $G\curvearrowright T^{\prime}$ is a splitting with a single orbit of edges, so it corresponds to either an amalgam $G=A*_{C}B$ or an HNN extension $G=A*_{C}$, with $C=G_{e}$ in either case. Let $e^{\prime}\in ET$ and $\mathfrak{h}^{\prime}\in\mathcal{H}(T^{\prime})$ be the images of $e$ and $\mathfrak{h}$ in $T^{\prime}$. If we choose orbit maps $f:G\to VT$ and $f^{\prime}:G\to VT^{\prime}$ so that $f^{\prime}$ is the composition of $f$ and the collapse map $T\to T^{\prime}$, then $f^{-1}(\mathfrak{h})=(f^{\prime})^{-1}(\mathfrak{h}^{\prime})$, so the halfspace for $G\curvearrowright T$ from Definition 2 corresponding to $\mathfrak{h}$ is identical to the halfspace for $G\curvearrowright T^{\prime}$ from Definition 2 corresponding to $\mathfrak{h}^{\prime}$ (assuming we work in the same Cayley graph of $G$ in both cases). Then by part (1) of the lemma, these halfspaces are quasi-isometric to either $\mathfrak{h}_{A},\mathfrak{h}_{B},\mathfrak{h}^{+}$ or $\mathfrak{h}^{-}$ from Definition 2.∎
Example 2.8.
Here is an example to show that halfspaces of group splittings are not necessarily quasi-isometrically embedded. Let $S$ be a closed surface with negative Euler characteristic and let $\phi:\pi_{1}(S)\to\pi_{1}(S)$ be an automorphism induced by a pseudo-Anosov homeomorphism of $S$. Let $G=\pi_{1}(S)\rtimes_{\phi}\mathbb{Z}$ be the semidirect product induced by $\phi$, and let $p:G\to\mathbb{Z}$ be the projection to the $\mathbb{Z}$ factor. Note that $G$ is a hyperbolic group [Thu22] (indeed it is the fundamental group of a closed hyperbolic 3-manifold). The map $p$ induces a splitting of $G$ (an action on $\mathbb{R}$), and the halfspaces correspond to the preimages $p^{-1}((-\infty,n])$ and $p^{-1}([n,\infty))$ for $n\in\mathbb{Z}$ (and their induced subgraphs in the Cayley graph of $G$). We claim that at least one of the halfspaces $p^{-1}((-\infty,0])$ and $p^{-1}([0,\infty))$ is not quasi-isometrically embedded in $G$. Indeed, suppose they are both quasi-isometrically embedded. Then for any $g_{1},g_{2}\in p^{-1}(0)$, there exist quasi-geodesics $\gamma^{-}$ and $\gamma^{+}$ (with respect to the metric on $G$) that lie in $p^{-1}((-\infty,0])$ and $p^{-1}([0,\infty))$ respectively and that connect $g_{1}$ to $g_{2}$. Moreover, the QI-constants of $\gamma^{-}$ and $\gamma^{+}$ are independent of $g_{1},g_{2}$. By the stability of quasi-geodesics in hyperbolic metric spaces, the quasi-geodesics $\gamma^{-}$ and $\gamma^{+}$ are uniformly close to each other, hence uniformly close to $p^{-1}(0)$. This implies that $p^{-1}(0)$ is a quasi-convex subgroup of $G$. But this is a contradiction, since an infinite normal subgroup of a hyperbolic group is quasiconvex if and only if it has finite index [MT94].
@@
As in Definition 2, let $p:X\to T$ be a tree of spaces for $G\curvearrowright T$. We must show that $p^{-1}(\mathfrak{h})$ is one-ended, so suppose for contradiction that $K\subseteq p^{-1}(\mathfrak{h})$ is a finite subgraph such that $p^{-1}(\mathfrak{h})-K$ has multiple unbounded components. We are assuming that $G$ and $G_{v}$ are one-ended, hence $X$ and $p^{-1}(v)$ are one-ended as well. In particular $p^{-1}(v)-K$ only has one unbounded component, call it $C_{1}$. Now consider the finite subgraph
@@
[…]
@@ 3 Pocsets
Elements $A,B\in\mathcal{P}$ are nested if $A\leq B$ or $B\leq A$. And $A,B\in\mathcal{P}$ are transverse if neither $A,B$ nor $A,B^{*}$ are nested. The width of $\mathcal{P}$ is the maximum number of pairwise transverse elements, if such a maximum exists, otherwise we say the width is $\infty$. We say that $\mathcal{P}$ is discrete if for any $A\leq B\in\mathcal{P}$ there are only finitely many $C$ with $A\leq C\leq B$.
@@
An ultrafilter $\omega$ is DCC (descending chain condition) if every infinite descending chain $A_{1}\geq A_{2}\geq A_{3}\geq\cdots$ terminates (i.e. there is $N$ with $A_{i}=A_{N}$ for all $i\geq N$).
@@
[Man20, Definition 9.10, Lemma 9.11 and Theorem 9.16] Let $\mathcal{P}$ be a pocset of finite width that admits at least one DCC ultrafilter. Then there is a CAT(0) cube complex $C=C(\mathcal{P})$, called the cubing of $\mathcal{P}$, such that:
@@
Proposition 3.2.
[Dun79, Theorem 2.1][Rol16, Theorem 9.1 and Proposition 9.3] Let $\mathcal{P}$ be a discrete pocset that contains no pair of transverse elements. Then $\mathcal{P}$ admits a DCC ultrafilter. Hence, by Proposition 3.1, $C(\mathcal{P})$ exists and is a tree.
Proof.
The result follows from the references given above, or it can easily be proved directly as follows. Choose $A\in\mathcal{P}$. Let $\omega=\{B\in\mathcal{P}\mid A\leq B\}\cup\{B\in\mathcal{P}\mid A^{*}\leq B\}$. Since $\mathcal{P}$ contains no pair of transverse elements, $\omega$ is an ultrafilter on $\mathcal{P}$. Furthermore, $\omega$ is DCC by the discreteness of $\mathcal{P}$. ∎
In this paper we will only deal with pocsets that satisfy Proposition 3.2. One source of pocsets will come from the following definition.
@@
[…]
@@ 4 Chopping up halfspaces
Remark 4.2.
One can transform the splitting $G\curvearrowright T^{\prime}$ back to $G\curvearrowright T$ by folding (in the sense of [BF91]). More precisely, there is a vertex $u_{0}^{\prime}\in VT^{\prime}$ and a family of edges $\mathcal{E}^{\prime}_{0}$ incident at $u_{0}^{\prime}$ such that folding together the edges in each family $g\mathcal{E}^{\prime}_{0}$ ($g\in G$) defines a $G$-equivariant map $T^{\prime}\to T$. Moreover, the image of $\mathcal{E}^{\prime}_{0}$ is the edge $e_{0}\in ET$ from (b), and the vertex $v_{0}\in VT$ from (d) is the endpoint of $e_{0}$ which is not the image of $u^{\prime}_{0}$. These facts can be deduced from the construction in Subsection 4.7.
@@
4.1 Strategy for Theorem 4.1
The rough strategy to prove Theorem 4.1 is as follows. We take a tree of spaces $p:X\to T$ associated to the splitting $G\curvearrowright T$ (Definition 2), and we take a halfspace $\mathfrak{h}_{0}\in\mathcal{H}(X)$ with more than one end. Say $\mathfrak{h}_{0}$ corresponds to an edge $e_{0}\in ET$. We let $\delta$ be some finite collection of edges which cuts $\mathfrak{h}_{0}$ into two unbounded components, and we chop up the halfspace $\mathfrak{h}_{0}$ using the $G_{e_{0}}$-translates of $\delta$. We define a tree $T_{e_{0}}$ which is somehow dual to this chopping, together with an action $G_{e_{0}}\curvearrowright T_{e_{0}}$. And then we do something similar to obtain the splitting $G\curvearrowright T^{\prime}$, but this time we chop up the entire space $X$ using all the $G$-translates of $\delta$ together with the existing walls $\hat{\mathfrak{h}}$ in $X$.
The difficulty comes in the details of how we define these trees $T_{e_{0}}$ and $T^{\prime}$. If we were only interested in defining $T_{e_{0}}$, then we could use the work of Dunwoody [Dun82, Theorem 1.1] to ensure that any two $G_{e_{0}}$-translates $g_{1}\delta,g_{2}\delta$ induce nested partitions of $\mathfrak{h}$, i.e. one of the complementary components of $g_{1}\delta$ is contained in one of the complementary components of $g_{2}\delta$, and from there we could define a dual tree $T_{e_{0}}$ whose edges correspond exactly to the $G_{e_{0}}$-translates of $\delta$ (using Section 3 or [Dun82, Theorem 4.1]). However, the situation becomes more complicated when we try to construct $T^{\prime}$. We want to chop up $X$ using the $G$-translates of $\delta$ together with the existing walls $\hat{\mathfrak{h}}$, but it is not clear how to produce a nested system of partitions of $X$ from this chopping. The solution is to choose the initial edge cut set $\delta$ in a very careful manner so that the $G_{e_{0}}$-translates of $\delta$ only chop up the halfspace $\mathfrak{h}_{0}$ and do not chop up any deeper halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ (Lemma 4.10). We are then able to construct a system of partitions of $X$ which is compatible with our chopping and which is “almost nested”. From this we define a pocset $\mathcal{P}$ whose elements are genuinely nested (Subsection 4.6), and we obtain the tree $T^{\prime}$ as the cubing of $\mathcal{P}$ (using Section 3 again). This construction relies on the edge cut set $\delta$ satisfying several properties, so it requires significant work to even construct a suitable $\delta$ (Subsections 4.2 and 4.3). In particular, in Subsection 4.2 we use a theorem of Krön [Krö10] (which is a refinement of Dunwoody’s theorem [Dun82, Theorem 1.1] mentioned above).
@@ 4.2 Minimal cuts
Definition 12. Let $Y$ be a connected simplicial graph. For $C\subseteq VY$, define the edge boundary of $C$, denoted $\delta C$, to be the set of edges in $Y$ with one vertex in $C$ and one vertex in VYCVY-C. A ray in $Y$ is an embedded one-way infinite path in $Y$. A cut is a subset $C\subseteq VY$ with $\delta C$ finite and such that the induced subgraphs of $C$ and VYCVY-C are both connected and contain rays. Note that, if $Y$ is locally finite, then a connected subgraph of $Y$ contains a ray if and only if it is unbounded. A cut $C$ is minimal if $|\delta C|$ is minimal.
@@
Clearly, $\mathcal{P}(Y,C)$ is a family of non-empty subsets of $VY$ that is closed under complementation. It remains to show that, for $x,y\in VY$, the set $S=\{A\in\mathcal{P}(Y,C)\mid x\in A,\,y\notin A\}$ is finite. Indeed, fix a path $\gamma$ from $x$ to $y$. Given $A\in S$, we either have $A=gC$ or A=VYgCA=VY-gC for some $g\in\operatorname{Aut}(Y)$. Now, $\delta(gC)$ determines the pair (gC,VYgC)(gC,VY-gC), so it suffices to show that there are only finitely many possibilities for $\delta(gC)$. But the path $\gamma$ contains an edge in $\delta(gC)$, and $\delta(gC)$ is an $\operatorname{Aut}(Y)$-translate of the finite edge set $\delta C$, so the local finiteness of $Y$ implies that there are only finitely many possibilities for $\delta(gC)$. ∎
@@
4.3 Halfspace cuts
We now turn to the proof of Theorem 4.1, which will be spread over Subsections 4.34.7. Let $G\curvearrowright T$ be a splitting as in the theorem, so a splitting with minimal action and finitely generated edge stabilizers, and with $G$ one-ended and finitely generated. And we assume that the splitting has some halfspace with more than one end. Let $p:X\to T$ be a tree of spaces associated to $G\curvearrowright T$ (Definition 2).
Define a halfspace cut to be a pair $(\mathfrak{h},C)$ where $\mathfrak{h}\in\mathcal{H}(X)$ and $C$ is a cut in $\mathfrak{h}$. As in Definition 4.2, let $\delta C$ denote the set of edges in $\mathfrak{h}$ with one vertex in $C$ and one vertex in $V\mathfrak{h}-C$. We are assuming that there exists a halfspace $\mathfrak{h}\in\mathcal{H}(X)$ with more than one end, so there exists $C\subseteq V\mathfrak{h}$ such that $(\mathfrak{h},C)$ is a halfspace cut.